OpenStudy (kayders1997):

Find integral of y/e^2y from 0 to 1

OpenStudy (kayders1997):

@Luigi0210 @ganeshie8

OpenStudy (kayders1997):

@welshfella @FortyTheRapper

OpenStudy (faiqraees):

What are your u and v ?

OpenStudy (kayders1997):

U=y v=idk

OpenStudy (kayders1997):

I'm trying to figure that out

OpenStudy (faiqraees):

v = 1/e^2y = e^-2y

OpenStudy (kayders1997):

That's Dv?

OpenStudy (kayders1997):

@FaiqRaees

OpenStudy (faiqraees):

No just v

OpenStudy (kayders1997):

Really?

OpenStudy (kayders1997):

Than my teacher taught us different

OpenStudy (faiqraees):

\[\large\rm v=\frac{1}{e^{2y}} = e^{-2y} \]

OpenStudy (faiqraees):

Btw you can also solve the question in the form 1/e^2y

OpenStudy (kayders1997):

Now don't I antififferentiate?

OpenStudy (faiqraees):

What's the formula you're using to integrate by parts?

OpenStudy (kayders1997):

You I gtg

OpenStudy (welshfella):

try integration by parts

OpenStudy (dannyo19):

Integration by parts: so you have two functions for your integrand \[I = y \times e ^{-2y}dy\]

OpenStudy (dannyo19):

Now you have for your integral I = u dv form

OpenStudy (dannyo19):

I simple u sub would work on this too.. Either way!